How many moles of NaOH are present in 18.0 mL of 0.230 M NaOH?
1
Expert's answer
2018-11-27T07:13:10-0500

Solution.

C = n/V n= CV

n(NaOH) = 0.018*0.23 =0.004 moles

Answer. 0.004 moles